OpenStudy (elonasushchik):

The equation of line WX is y = -2x - 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (-1, -2).

OpenStudy (elonasushchik):

Y = ½ x + 3/2 Y = -1/2 x + 3/2 Y = ½ x - 3/2 Y = 1/-2 x - 3/2

OpenStudy (elonasushchik):

Those are the choices^^^^^^

OpenStudy (elonasushchik):

@AlexandervonHumboldt2

OpenStudy (elonasushchik):

@e.mccormick

OpenStudy (alexandervonhumboldt2):

yeah so perpendivular lines have negative reciprocla slopes

OpenStudy (elonasushchik):

okay soo... I think its the second choice.

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (elonasushchik):

i was thinking 2 or 3rd.. but thanx anyways @AlexandervonHumboldt2 <3!!!